OpenStudy (anonymous):

Identify the correct slope and y-intercept of the equation 2x + 4y = 8 slope = -one half; y-intercept = (0, 2) slope = -2; y-intercept = (0, 4) slope = -2; y-intercept = (0, 8) slope = -one half; y-intercept = (0, 8)

OpenStudy (anonymous):

2x+4y=8 4y=-2x+8 on compating with y=mx+c; slope(m) =-2 and y intecept is c=(0,8) therefore option 3 is your suitable answer...

OpenStudy (anonymous):

slope will be -1/2 and intercept (0,2)..

OpenStudy (anonymous):

uh @t_soulb you are absolutely wrong..mine is the correct answer..@tristynhart go through my answer its 100% correct..

OpenStudy (anonymous):

1st reduce coefficient of y to one then you will get the 'c 'intercept ....

OpenStudy (anonymous):

you comparing with y = mx+c , so 4y =-2x+8 will reduce to y = -x/2 + 2 then compare dear..

OpenStudy (anonymous):

ya sorry i forget to reduce it..ya you are right..first reduce it and compare y = -x/2 + 2 and compare with y=mx+c we get the answer -1/2 and (0,2) so sorry for my first answer..this one is correct...

OpenStudy (anonymous):

no problem ..